# Are Helium Atmospheres Hiding as Volatile-rich in Current Exoplanet Analysis Frameworks?

### v1

@@ Abstract
The increasing number of detailed exoplanet observations offers an opportunity to refine our analyses and interpretations, and uncover previously-hidden trends. One such trend is unexpectedly high water abundances (110%) for several hot Jupiters. Here, we show that such atmospheres may instead be helium-rich, and the perceived high water abundances simply biases in analyses compensating for “ad hocly” fixed helium-to-hydrogen ratio (He/H2 = 0.157). As transmission spectra constrain the atmospheric scale height (), a helium-enriched atmosphere can be misinterpreted as H2-dominated water-rich to bring the mean molecular weight () to intermediate values (415) when He/H2 is fixed. We present a proof-of-concept reanalysis of HD 209458 b’s JWST transmission spectrum treating He/H2 as a free parameter, resulting in sets of solutions with helium enrichment and non-extreme water abundance. The He-enhancement found requires an order of magnitude less enrichment/evolution than does a water abundance of 10%, suggesting that the He-rich hypothesis is a viable alternative. We argue that He enhancement must be considered to reliably constrain atmospheric composition, be sensitive to a more diverse planetary population, and ultimately yield robust metallicity trends to inform formation/evolution pathways. Looking ahead, we suggest leveraging insights from differences in pressure-broadening effects, outflow measurements, and atmospheric chemistry to help disentangle reliably between He- and volatile-rich atmospheres—while recognizing that existing opacity/outflow/chemistry models may need targeted upgrades to reach the required fidelity level.
1 Introduction
@@
The field of exoplanetary science has undergone a remarkable transformation in the last decades, evolving from the detection of a handful of exoplanets to the detailed characterization of exoatmospheres. Yielding observations with unprecedented information content, JWST is at the forefront of this ongoing transformation.
@@
As the community learns to leverage and interpret the novel data, fresh challenges arise. Reliably translating a planet’s transmission spectrum into physical constraints now requires exquisite knowledge of its host’s heterogeneities (e.g., Lim et al., 2023; Moran et al., 2023; Rackham & de Wit, 2023; Berardo et al., 2024; TRAPPIST-1 JWST Community Initiative et al., 2024), and of light-matter interactions (e.g., Fortney et al., 2019; Greaves et al., 2021; Ranjan et al., 2020; Niraula et al., 2022). Beyond model upgrades, it is also key to keep (re)assessing the complexity we allow for in our models to ensure they reach a sufficient fidelity. For example, using simple 1D models can prevent the identification of day-night temperature asymmetries and lead to underestimated temperatures by up to 1,000K (MacDonald et al., 2020).
@@
Here, we examine the impact of a common assumption behind the interpretation of giant exoplanets’ spectra: the helium-to-hydrogen (He/H2) ratio identical to that of Jupiter (He/H2 = 0.157, Conrath et al., 1987). We choose to challenge this assumption considering the unexpectedly high water abundances () reported for a handful of giants (e.g., HD 209458 b, WASP-39 b, and WASP-80 b, Niraula et al., 2023; Bell et al., 2023; Xue et al., 2024), which are 2 orders of magnitude higher than those found in the solar system (Asplund et al., 2009; Atreya et al., 2016). As most JWST datasets target giant exoplanets under irradiations beyond those found in the solar system, freeing atmospheric retrieval codes from an assumption anchored in solar-system observations regarding He/H2 seems well motivated.
In fact, Hu et al. (2015) first presented the hypothesis of a He-dominated atmosphere for a Neptune-sized planet (GJ 436 b possibly turned He-rich via atmospheric hydrodynamic escape). Since other studies have investigated a range of fractionation regimes for similar and smaller planets (e.g., Malsky et al., 2023; Cherubim et al., 2024; Lammer et al., 2025).
2 Concept
@@
A planet’s transmission spectrum can yield independent constraints on its temperature, composition, and pressure profile (de Wit & Seager, 2013), including its atmospheric scale height ()111k is the Boltzmann constant, T is the temperature, is the molecular weight, and is the gravity.. This can result in tight constraints on the atmospheric mean molecular weight (e.g., Niraula et al., 2025), notably for planets with independent mass constraints from radial velocities or transit timing variations. Therefore, if deviates from the default 2.3 due to helium enrichment (or the presence of a dominant atmospheric species without strong absorption features–i.e., not accounted/fitted for in a retrieval), retrieval codes tend to artificially amplify the abundance of the dominant absorber (often water).
@@
We thus hypothesize that the high water content (or, in fact, metallicity) reported recently in the literature may result from biases stemming from assuming He/H2 = 0.157. This forces the dominant absorbers with saturated absorption features to higher abundances to tweak and match the value recorded by the transmission spectrum. This would mean that current (fixed-He/H2) retrieval frameworks are blind to He-rich atmospheres and risk yielding artificially enhanced metallicity trends.
3 Proof of Concept
To test this idea, we reanalyzed HD 209458 b’s JWST spectrum while fitting for He/H2. We chose HD 209458 b given the high water abundance reported in the literature (O(10)%, see Xue et al., 2024).
3.1 Data
@@
The spectrum we used in our analysis is derived from two 8.01-hour transit observations with JWST-NIRCam (GTO 1274; PI: Lunine) and was reduced using Eureka (Bell et al., 2022). We note that each of the reported uncertainties were inflated by a factor of 1.4 to account for underestimated noise, which explains the apparent lack of large residuals in the model fits to the data (see Figure 2).
3.2 Analysis Framework
@@
To perform our atmospheric retrievals, we use tierra222https://github.com/disruptiveplanets/tierra (Niraula et al., 2022, 2023, 2025). Unlike many previous retrievals that fix the helium-to-hydrogen ratio (He/H2) to a canonical value of 0.157, we treat He/H2 as a free parameter to investigate its impact on the retrieved chemical abundances—and test for helium enrichment or depletion.
@@
For our retrievals, we consider eight different molecules: , , , , , , , and , applying a uniform prior on the log mixing ratio ranging from -0.5 to -12, and check later that none of the PPDs are affected by this choice (i.e., the retrieved molecular abundances did not pile up at the upper bounds of the prior, as has been seen in some previous retrievals). We also vary the planetary mass, placing Gaussian priors based on the published radial velocity mass (Bonomo et al., 2017).
3.3 Results
@@
Our retrievals return a broad range of possible values for He/H2 and a clear negative trend between He/H2 and the H2O abundance (Figure 2). As anticipated (Section 2), the increasing He/H2 ratio also lessens the reliance on H2O abundance in increasing , which results in H2O’s abundance being less constrained then.
3.3.1 A Perspective from Molecular Abundances
@@
To further highlight this, we split our ensemble of solutions into those with Log10[He/H2] and in Figure 2. This helps show that retrievals with higher He/H2 ratios consistently yield lower water abundances, while lower He/H2 ratios constrain water abundance to higher values and a much tighter part of the parameter space (Figure 2, middle panel). Specifically, we find a water abundance of for the Log10[He/H2] sample and for Log10[He/H2].
@@
Conversely, the abundance of a non-dominant absorber like H2S is affected marginally by higher He/H2 ratios. This is because H2S contributes negligibly to compared to more abundant species with saturated absorption features like H2O. As a result, when the He/H2 ratio increases and rises, the retrieval framework does not need to significantly adjust/compensate the inferred H2S abundance to maintain a good fit.
3.3.2 A Perspective from the Mean Molecular Weight
@@
We can also look at this matter through the lens of the atmospheric mean molecular weight. Transmission spectra primarily constrain through their sensitivity to the atmospheric scale height (de Wit & Seager, 2013). In our retrievals, we independently computed by summing the retrieved mixing ratios of all species, weighted by their molecular masses. The peak of the distribution lies significantly beyond the canonical value of expected for a Jupiter-like composition (Figure 3), suggesting a substantially-heavier atmosphere. If one instead assumes a fixed , the retrieval must compensate by artificially inflating the abundance of the dominant absorber (typically water). This degeneracy highlights how fixed assumptions about He/H2 can directly lead to overestimates of atmospheric water content (or metallicity).
4 Discussion
While the JWST era has only just begun, many giant exoplanets have already been found to exhibit surprisingly-high water abundances. For example—in addition to HD 209458 b (see previous Sections and Xue et al., 2024), Niraula et al. (2023) retrieved a water abundance of () for WASP-39 b, Bell et al. (2023) reported () for WASP-80b, and Mayo et al. (2024) found a high water abundance of () for WASP-166 b. These values are 2 orders of magnitude higher than those found in the solar system333In fact, these abundances are often underestimated because many studies present estimates from posterior probability distributions clearly bouncing against (i.e., truncated by) the upper bounds priors. For example, while Bell et al. (2023) reports a () their PPD hits the upper bound well within 1 (see their Extended Data Fig. 3) meaning that without such hard edge the favored water abundance would likely be closer to ..
In contrast, He-rich atmospheres require 1 order of magnitude enrichment (or escape). While He-rich giants may sound exotic at first, considering the aforementioned data- and enrichment-driven arguments, we pursue and explore possible origins for and means to disentangle between He-rich and H2-dominated volatile-rich atmospheres in this Section.
4.1 Giant exoplanets with Helium atmospheres?
@@
While hydrogen is the most abundant element in gas giants in our solar system, it is conceivable that a giant exoplanet becomes helium-rich—e.g., under conditions where hydrogen is more readily photoevaporated. Figure 4 illustrates a schematic “before and after” evolution for a giant atmosphere having undergone such an enrichment. In the initial state (left), the planet has a stratified structure with a molecular hydrogen-dominated, helium-poor upper atmosphere, a helium rain layer where helium droplets condense and sink. Over time, as hydrogen is selectively lost from the upper layers via photoevaporation (Owen & Wu, 2013, 2017), the planet transitions to the state shown on the right: a helium-rich atmosphere where the helium-enriched regions dominate the transmission spectroscopy signal with an intermediate and thus a lower . This mechanism provides a natural explanation for why some exoplanets, especially Neptune-sized and smaller (e.g., Hu et al., 2015; Malsky et al., 2023; Cherubim et al., 2024; Lammer et al., 2025), might exhibit atmospheric compositions that are more He-rich than those of Jupiter or Saturn.
4.2 Other Possible Explanations
@@
First and foremost, these worlds may very well be truly water-rich (see set of blue solutions in Figure 2). Although such high abundances are not anticipated based on solar system giants or equilibrium chemistry models (e.g., Mollière et al., 2017), missing physics and chemistry in current formation or atmospheric evolution models could allow for heavier-than-expected compositions in highly irradiated environments.
@@
4.3 He-rich? Volatile-rich? Which is Which?
While JWST offers unparalleled precision for constraining atmospheric compositions, directly constraining the He/H2 ratio for giant exoplanets is challenging (e.g., Fig. 2). We discuss here three sources of complementary information to help disentangle between He-rich and volatile-rich (i.e., H2-dominated) exoplanets.
4.3.1 Scattering, Absorbing, and Broadening Behaviors
@@
To highlight these three spectral effects and showcase how they may help break the degeneracy between He-rich and (H2-dominated) water-rich atmospheres, we present in the left panel of Figure 5 the extended best-fit models to the HD 209458 b spectrum presented in Figure 2. Rayleigh scattering values are taken from Wilmouth & Sayres (2019) and Sneep & Ubachs (2005) for helium and hydrogen, respectively. It shows that the best fit models deviate substantially from each other in the VIS/NIR, primarily due to He’s scattering being 2 orders of magnitude smaller than H2’s. In clear atmospheres, it would thus be possible to identify the differences between two scattering signatures. Unfortunately, close-in planets often have cloudy/hazy atmospheres supported by enhanced photochemical production (see HST data for HD 209458 b in Figure 5 and, e.g., Sing et al., 2016). We thus expect that disentangling between a helium- and hydrogen-dominated atmosphere solely from the scattering signal to be challenging.
@@
The next clear difference between both gases relates to the presence/absence of H2 CIA features. We point in particular to the 2.4m feature able to peak between water and methane features (Figure 5, left panel). Relative measurements of out-of-feature transit depths (i.e., 2.4m vs 1.7 and 3.4 m) can provide direct insights into the contribution of H2’s CIA and leverage part of transmission spectra less sensitive to the signals that can overpower scattering slopes (e.g., hazes and stellar contamination, see Triaud et al., 2024). Unfortunately, the aforementioned out-of-feature windows will close for volatile-rich atmospheres (Figure 5, left panel), preventing the detection H2’s CIA.
@@
Finally, we turn to the secondary spectroscopic effect associated the pressure-broadening of other species’ lines. Niraula et al. (2022) demonstrated that information (or the lack thereof) associated with pressure broadening in central to an exoplanet transmission spectrum, while Wiesenfeld et al. (2025) reported that JWST-quality spectra can be sensitive to difference in the pressure-broadening coefficients down to 10%. Fortunately, the difference between the pressure-broadening coefficients of most gases by H2 vs by He is one order of magnitude above that sensitivity limit (e.g., Tan et al., 2022, and references therein). We showcase in the right panel of Figure 5 the detectable differences in the transmission spectrum of an exoplanet for which only the pressure-broadening coefficients have been changed; 50-ppm level across the spectrum, esp. in the out-of-feature regions (more pressure-broadening sensitive). As retrievals can hide such a signal, reliably sensing and interpreting such H2-vs-He spectral differences will require targeted upgrades of opacity models (Tan et al., 2022; Niraula et al., 2022; Wiesenfeld et al., 2025).
4.3.2 Outflow Fractionation

### v2

# Helium Atmospheres May Hide in Current Exoplanet Analysis Frameworks

@@ Abstract
The increasing number of detailed exoplanet observations offers an opportunity to refine our analyses and interpretations. Here, we show that atmospheres that appear volatile-rich and/or cloudy may instead be helium-rich. As transmission spectra constrain the atmospheric scale height (), a He-enriched atmosphere can be misinterpreted as H2-dominated water-rich to bring the mean molecular weight () to intermediate values (410) when He/H2 is fixed. Similarly, a cloud deck can reduce the spectral features, and thus the apparent (i.e., cloud-free equivalent) . We present a proof-of-concept reanalysis of HD 209458 b’s JWST transmission spectrum treating He/H2 as a free parameter, resulting in sets of He-rich solutions. We argue that He enhancement must be considered to reliably constrain atmospheric composition, be sensitive to a more diverse planetary population, and ultimately yield robust trends to inform formation and evolution pathways. Looking ahead, we suggest leveraging insights from differences in pressure-broadening effects, outflow measurements, and atmospheric chemistry to disentangle reliably between He-, volatile-rich, and cloudy atmospheres—while recognizing that associated models need targeted upgrades to reach the fidelity level required to this end.
I Introduction
@@
The field of exoplanetary science has undergone a remarkable transformation in the last decades, evolving from the detection of a handful of exoplanets to the increasingly-detailed characterization of exoatmospheres. Yielding data with unprecedented information content, JWST is at the forefront of this transformation.
@@
As the community learns to leverage and interpret the novel data, fresh challenges arise. Reliably translating a planet’s transmission spectrum into physical constraints now requires exquisite knowledge of its host’s heterogeneities (e.g., Lim2023; Moran2023; Rackham2023; Berardo2024; TJCI2024), and of light-matter interactions (e.g., Fortney2019; Greaves2021; Ranjan2020; Niraula2022). Beyond model upgrades, it is also key to keep (re)assessing the complexity we allow for in our models to ensure they reach a sufficient fidelity. For example, using simple 1D models can prevent the identification of day-night temperature asymmetries and lead to underestimated temperatures by up to 1,000K (MacDonald2020).
@@
Here, we examine the impact of a common assumption behind the interpretation of giant exoplanets’ spectra: a helium-to-hydrogen (He/H2) ratio identical to that of Jupiter (He/H2 = 0.157, Conrath1987). As most JWST datasets target giant exoplanets under irradiations beyond those found in the solar system, freeing atmospheric retrieval codes from an assumption anchored in solar-system observations regarding He/H2 seems well motivated. Hu2015 first presented the hypothesis of a He-dominated atmosphere for a Neptune-sized planet (GJ 436 b possibly turned He-rich via atmospheric hydrodynamic escape). Since other studies have investigated a range of fractionation regimes for similar and smaller planets (e.g., Malsky2023; Cherubim2024; Lammer2025).
II Concept
@@
A planet’s transmission spectrum can yield independent constraints on its temperature, composition, and pressure profile (deWit2013), including its atmospheric scale height (, with Boltzmann’s constant, the temperature, the molecular weight, and the gravity—see also Miller2009). This can result in tight constraints on the atmospheric mean molecular weight (e.g., Niraula2025; dewit2025), notably for planets with independent mass constraints from radial velocities or transit timing variations. Therefore, if deviates from the default 2.3 due to helium enrichment (or the presence of a dominant atmospheric species without strong absorption features–i.e., not accounted/fitted for in a retrieval), retrieval codes artificially amplify the abundance of the dominant absorber (often water)—or involve clouds (see Section IV.2 and, e.g., Berta2012).
@@
We thus hypothesize that high water content (or, in fact, metallicity) may result from biases stemming from assuming He/H2 = 0.157. This forces the dominant absorbers with saturated absorption features to higher abundances to tweak and match the value recorded by the transmission spectrum. This would mean that current (fixed-He/H2) retrieval frameworks are blind to He-rich atmospheres and risk yielding artificially enhanced metallicity trends. We discuss in Section IV.2 that the same goes for clouds.
III Proof of Concept
III.1 Data
@@
To test this idea, we reanalyze HD 209458 b’s JWST spectrum while fitting for He/H2. The spectrum used is derived from two 8.01-hour transit observations with JWST-NIRCam (GTO 1274; PI: Lunine) and was reduced using Eureka (Bell2022). We note that each of the reported uncertainties were inflated by a factor of 1.4 to account for underestimated noise, which explains the apparent lack of large residuals in the model fits to the data (see Figure 2).
III.2 Analysis Framework
@@
To perform our atmospheric retrievals, we use tierra111https://github.com/disruptiveplanets/tierra (Niraula2022; Niraula2023; Niraula2025). Unlike many previous retrievals, we treat He/H2 as a free parameter to assess its impact on the retrieved abundances, and test for He enrichment or depletion. For this first proof-of-concept application, we assume the atmosphere to be cloud-free.
@@
For our retrievals, we consider eight different molecules: , , , , , , , and , applying a uniform prior on the log mixing ratio ranging from -0.5 to -12, and check later that none of the PPDs are affected by this choice (i.e., the retrieved molecular abundances did not pile up at the upper bounds of the prior, see Section IV.2 ). We also vary the planetary mass, placing Gaussian priors based on the published radial velocity mass (Bonomo2017).
III.3 Results
@@
Our cloud-free retrievals return a broad range of possible values for He/H2 and a clear negative trend between He/H2 and the H2O abundance (Figure 2). As anticipated (Section II), the increasing He/H2 ratio also lessens the reliance on H2O abundance in increasing , which results in H2O’s abundance being less constrained then.
III.3.1 A Perspective from Molecular Abundances
@@
To highlight this further, we split our ensemble of solutions into those with Log10[He/H2] and in Figure 2. This helps show that retrievals with higher He/H2 ratios consistently yield lower water abundances, while lower He/H2 ratios constrain water abundance to higher values and a much tighter part of the parameter space (Figure 2, middle panel). Specifically, we find a water abundance of for the Log10[He/H2] sample and for Log10[He/H2].
@@
Conversely, the abundance of a non-dominant absorber like H2S is affected marginally by higher He/H2 ratios. H2S contributes negligibly to compared to more abundant species with saturated absorption features like H2O. As a result, when the He/H2 ratio increases and rises, the retrieval framework does not need to significantly adjust/compensate the inferred H2S abundance to maintain a good fit.
III.3.2 A Perspective from the Mean Molecular Weight
@@
We can also look at this matter through the lens of . Transmission spectra primarily constrain through their sensitivity to (Miller2009; deWit2013). In our retrievals, we independently computed by summing the retrieved mixing ratios of all species, weighted by their molecular masses. The peak of the distribution lies significantly beyond the canonical value of expected for a Jupiter-like composition (Figure 3), suggesting a substantially-heavier atmosphere. If one instead assumes a fixed , the retrieval must compensate by artificially inflating the abundance of the dominant absorber (typically water). This degeneracy highlights how fixed assumptions about He/H2 can directly lead to overestimates of atmospheric water content (or metallicity).
IV Discussion
IV.1 Giant exoplanets with Helium atmospheres?
While He-rich giants may sound exotic at first, He-rich atmospheres require 1 order of magnitude enrichment compared to the gas giants of the solar system. This is over an order of magnitude less enrichment than needed for a volatile-rich giants (e.g., water). Although hydrogen is the most abundant element in solar-system gas giants, it is conceivable that a giant exoplanet becomes He-rich—e.g., under conditions where hydrogen is more readily photoevaporated.
@@ IV.1 Giant exoplanets with Helium atmospheres?
Figure 4 illustrates a schematic “before and after” evolution for a giant atmosphere having undergone such an enrichment. In the pre-migration initial state (left), the planet has a stratified structure with a molecular hydrogen-dominated, helium-poor upper atmosphere, a helium rain layer where helium droplets condense and sink. Over time, as hydrogen is selectively lost from the upper layers via photoevaporation (Owen2013; Owen2017), the planet transitions to the state shown on the right: a helium-rich atmosphere where the helium-enriched regions dominate the transmission spectroscopy signal with an intermediate and thus a lower . This mechanism may explain why some exoplanets, especially Neptune-sized and smaller (e.g., Hu2015; Malsky2023; Cherubim2024; Lammer2025), might exhibit atmospheres more He-rich than Jupiter’s or Saturn’s.
While an atmosphere-only helium enrichment (or inverted metallicity gradients) may at first be predicted as unstable (e.g., Thorngren2019), new mechanisms have been introduced to explain the newly-discovered inverted metallicity gradient in Jupiter (Muller2024)—e.g., a stronger subadiabaticity to stabilize the outer stable layer and support, notably an inverted helium gradient (Nettelmann2025). The alternative of a (quasi-)global helium enrichment of the planet presents the same challenge at first: interior models predicts Jupiter-mass objects composed primarily of helium to be as small as (e.g., Stevenson1976). However, the same models predicts radii up to 2x smaller than observed for a subpopulation of hot Jupiters (a phenomenon coined “the mystery of the inflated radii”, see Thorngren2018, and references therein). In other words; to avoid biases and ensure that we fully harvest the potential of this data-rich era of astronomy, it is important to appreciate that—while helpful—models are still subject to observation-driven upgrades, as recently highlighted by the discovery of a helium-rich (and carbon-rich) “planet” (Zhang2025).
IV.2 Other Possible Explanations
@@
First and foremost, these worlds may very well be truly water-rich (see set of blue solutions in Figure 2). Although such high abundances are not anticipated based on solar system giants or equilibrium chemistry models (e.g., Molliere2017), missing physics and chemistry in current formation or atmospheric evolution models could allow for heavier-than-expected compositions in highly irradiated environments.
@@
Third, reduced spectral features can also result from the presence of high-altitude clouds (Berta2012). To showcase this additional dimension of relevance, we relax the could-free assumption in our analysis of HD 209458 b’s spectrum. Doing so, we find that for cloudy scenarios (i.e., with cloud top above 100 mbar) is not constrained to intermediate values, but can be consistent with conventional He/H2 ratios (Fig. 6).
IV.3 He-rich? Volatile-rich? Which is which?
While JWST offers unparalleled precision for constraining atmospheric compositions, directly constraining the He/H2 ratio for giant exoplanets is challenging. We discuss here three sources of complementary information to help disentangle between the He- vs volatile-rich scenarios—see Benneke2013 for further details on unambiguously identifying cloud signatures.
IV.3.1 Scattering, Absorbing, and Broadening Behaviors
@@
To highlight these three spectral effects and showcase how they may help break the degeneracy between He-rich and (H2-dominated) water-rich atmospheres, we present in the left panel of Figure 5 the extended best-fit models to the HD 209458 b spectrum presented in Figure 2. Rayleigh scattering values are taken from wilmouth2019 and sneep2005 for helium and hydrogen, respectively. It shows that the best fit models deviate substantially from each other in the VIS/NIR, primarily due to He’s scattering being 2 orders of magnitude smaller than H2’s. In clear atmospheres, it would thus be possible to identify the differences between two scattering signatures. Unfortunately, close-in planets often have cloudy/hazy atmospheres supported by enhanced photochemical production (see HST data for HD 209458 b in Figure 5 and, e.g., Sing2016). We thus expect that disentangling between a helium- and hydrogen-dominated atmosphere solely from the scattering signal to be challenging.
@@
The next clear difference between both gases relates to the presence/absence of H2 CIA features. We point in particular to the 2.4m feature able to peak between water and methane features (Figure 5, left panel). Relative measurements of out-of-feature transit depths (i.e., 2.4m vs 1.7 and 3.4 m) can provide direct insights into the contribution of H2’s CIA and leverage part of transmission spectra less sensitive to the signals that can overpower scattering slopes (e.g., hazes and stellar contamination, see Triaud2024). Unfortunately, the aforementioned out-of-feature windows will close for volatile-rich atmospheres (Figure 5, left panel), preventing the detection H2’s CIA.
@@
Finally, we turn to the secondary spectroscopic effect associated the pressure-broadening of other species’ lines. Niraula2022 demonstrated that information (or the lack thereof) associated with pressure broadening in central to an exoplanet transmission spectrum, while Wiesenfeld2025 reported that JWST-quality spectra can be sensitive to difference in the pressure-broadening coefficients down to 10%. Fortunately, the difference between the pressure-broadening coefficients of most gases by H2 vs by He is one order of magnitude above that sensitivity limit (e.g., Tan2022, and references therein). We showcase in the right panel of Figure 5 the detectable differences in the transmission spectrum of an exoplanet for which only the pressure-broadening coefficients have been changed; 50-ppm level across the spectrum, esp. in the out-of-feature regions (more pressure-broadening sensitive). As retrievals can hide such a signal, reliably sensing and interpreting such H2-vs-He spectral differences will require targeted upgrades of opacity models (Tan2022; Niraula2022; Wiesenfeld2025).
IV.3.2 Outflow Fractionation